# From Text to Emoji: How PEFT-Driven Personality Manipulation Unleashes the Emoji Potential in LLMs

### v2

@@ 2 Methodology
Personality Manipulation Methods and Metrics
The In-Context Knowledge Editing (IKE) method (Zheng et al. (2023)) was used as a baseline to manipulate embedded knowledge in LLMs, serving as a comparative foundation for evaluating prompt-based versus fine-tuning techniques in personality manipulation. The same prompt from Mao et al. (2023), provided in A.6, ensured consistency in comparing IKE with PEFT. We excluded methods like MEND due to inconsistent and gibberish outputs. Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLoRA), was selected to reduce computational cost while maintaining performance. The process began by preparing the Personality dataset, formatting an additional column as <s>[INST] Question [/INST] Answer </s>. Models were loaded from Hugging Face using 4-bit quantization via BitsAndBytes (Dettmers et al. (2023b)) for efficient processing and storage. The temperature was set to 1. The hyperparameters, chosen for their balance of efficiency and performance (Houlsby et al. (2019); Dettmers et al. (2024)), are detailed in A.4. After training, fine-tuned LoRa weights were merged with the original model for evaluation.
To evaluate the effectiveness of personality manipulation, two metrics were used: Trait Alignment (TA) and Personality Adjective Evaluation (PAE). TA is computed using the metric classifier by comparing predicted labels $\hat{y}_{i}$ with true labels $y_{i}$, which correspond to the original target personality traits from the dataset. For a dataset with $N$ instances, TA is given by $\text{TA}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}(\hat{y}_{i}=y_{i})$, where $\mathbb{1}(\hat{y}_{i}=y_{i})$ is 1 if the prediction matches the true label and 0 otherwise, providing an average alignment between predictions and the intended personality traits. PAE, inspired by Mao et al. (2023), uses Chain of Thought (CoT) prompting (Wei et al. (2022)), where a larger LLM (here GPT-4 (Brown (2020))) scores generated text on a 1-5 scale based on its alignment with the target trait. The PAE score is calculated as the difference between the score of generated text and original text from the dataset, with higher difference indicating that the generated text after PEFT more effectively captures the intended personality traits compared to the original text. The final PAE is the mean of these score differences across all instances, $\text{PAE}=\frac{1}{N}\sum_{i=1}^{N}s_{i}$, where $s_{i}$ is the score difference for each instance $i$. Details of the prompt are provided in A.5.
3 Result and Discussion
The performance of PEFT and IKE is presented in Figure 4 and detailed in Table 11 in A.9. PEFT consistently outperforms IKE in TA across most personality traits and models, indicating more reliable manipulation. Although IKE occasionally exceeds PEFT in PAE for certain traits by capturing nuanced text alignment, PEFT provides deeper and more stable personality manipulation and better scalability when applied across multiple traits. We additionally employed prompting and manual verification to validate effective manipulation across all traits (in A.8). During inference, fine-tuned Llama-2-7B-Chat and Mistral-7B-Instruct models generated emojis, despite no emojis being present in the PEFT training data. In contrast, the original models produced generic responses without emojis for the same inputs. This behaviour is likely due to pre-training on diverse corpora containing emoji patterns, with PEFT amplifying these latent tendencies. To test this, we performed neuron activation analysis, a mechanistic interpretability method (Bereska and Gavves (2024)), finding that specific neurons in Mistral-7B-Instruct and Llama-2-7B-Chat were responsible for emoji generation, becoming more active during conversational prompts. This suggests pre-training on informal data influenced these behaviours. In contrast, Llama-3-8B-Instruct showed no such activation, implying that these tendencies were either not learned or suppressed during fine-tuning when first developed (see A.7). To verify that the emojis were intentional and not random artefacts, we also conducted an in-context learning (ICL) explainability analysis through prompting, confirming that emoji usage aligned with intended personality traits. This analysis followed the same method as manipulation result validation (in A.8) but focused specifically on emojis. The Emoji-to-Sentence Ratio (ESR) was used to measure the frequency of emojis in generated responses after PEFT, calculated as ESR=Sentences with emojisTotal $\text{ESR}=\frac{\text{Sentences with emojis}}{\text{Total sentences}}$. Table 1 shows the results from mechanistic interpretability and ICL analysis, highlighting frequently used emojis and their corresponding ESRs. Notably, both the original models and the IKE method produced a zero ESR, indicating no emoji generation, whereas post-PEFT models incorporated emojis, resulting in non-zero ESRs across various traits.
Model Method Trait ESR ICL (Emoji) Mechanistic (Neuron)
Mistral-7B-Instruct Original All Traits 0 - Distributed activation
suggests the model
handles emoji generation
across multiple neurons,
indicating greater flexibility.
@@ 3 Result and Discussion
Llama-2-7B-chat Original All Traits 0 - Single peak activation,
suggests the model has specialised neurons for
recognising and processing emoji generation.
@@
Table 1: ESR, ICL Emoji, and Neuron Activation in Generated Text for Different Personality Traits
@@
Llama-2-7B-Chat predominantly produced emojis for Extraversion and Neuroticism, with Extraversion showing the highest emoji-to-sentence ratio at 0.995, where nearly every sentence included an emoji. ICL emojis for Extraversion were positive, such as [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image], reflecting Extraversion qualities. Neuroticism had a ratio of 0.255, using more negative emojis like [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]. Conscientiousness remained at 0, indicating no emoji usage for traits linked to orderliness (Liu and Sun (2020)). Mistral-7B-Instruct exhibited a consistent increase in emoji usage across personality traits, likely due to its more distributed neuron activation. This was especially pronounced for Openness (0.925) and Conscientiousness (0.82), where creative and productive emojis like [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image] were frequently used. Neuroticism had a moderate ratio (0.575), featuring negative emojis such as [Uncaptioned image] and [Uncaptioned image]. Thus, the PEFT results align with personality traits, enhancing the LLMs’ expressive and contextually appropriate content (Kennison et al. (2024)).
4 Limitations and Future Work
This study assesses personality manipulation in LLMs using the Big Five model, revealing inconsistencies such as high TA in Neuroticism (0.975) and low TA in Agreeableness (0.065) in Llama-2-7B-Chat. The disparity may stem from Neuroticism’s emotional salience versus Agreeableness’s context dependence. Future work should enhance fine-tuning consistency and consider continuous personality measures to capture nuances missed by classifiers. Additionally, exploring alternative psychometric models like 16PF (Cattell and Mead (2008)) could provide further insights into personality manipulation. Future work could also explore how model size correlates with manipulation effectiveness and whether there is a threshold where scaling yields diminishing returns.
Appendix A Appendix / supplemental material
@@ A.1 Personality Manipulation Dataset
Personality Manipulation dataset consists of $5000$ instances, with $4000$ allocated for training and $1000$ for testing. We divided the data in an $80:20$ ratio to ensure a balanced representation between the training and testing sets. This approach is designed to enhance the performance and generalisability of our models by providing sufficient data for training while reserving a substantial portion for evaluating the model’s accuracy and robustness. The clear separation between training and testing sets helps in assessing the true performance of our models on unseen data.
@@
Instruction: Exhibit the trait of Target Personality when answering the question to express opinion on the certain Edit Topic
@@
Table 2: Prompt used for Few-shot Prompting in Dataset Generation
@@
For instance, as visible in the figure 6, the word cloud for extraversion showcases words that reflect the energetic, outgoing, and sociable nature of individuals with this personality trait. Prominent terms such as "absolutely," "always," "believe," "love," and "thrilling" emphasise the enthusiasm, positivity, and social engagement typical of extraverts. Similarly, the word cloud for neuroticism reveals words associated with emotional instability, anxiety, and self-consciousness. Key terms like "feel," "worry," "sometimes," "make," and "bit" are prominently displayed, reflecting the frequent emotional fluctuations and concerns typical of neurotic individuals. Words like "anxious," "overwhelming," and "struggle" further emphasise the challenges faced by those with high levels of neuroticism. Collectively, these word clouds provide a rich visual representation of the lexical diversity associated with each personality trait, offering valuable insights into the distinct characteristics and behavioural tendencies that define them.
@@ A.1.1 Text Analysis
In conclusion, LDA-based topic modeling reveals significant insights into how different themes resonate with various personality traits. This understanding is crucial for the manipulation of personalities in language models, allowing for the creation of content that is tailored to engage specific personality profiles effectively. By leveraging these insights, content can be crafted to not only align with individual preferences but also to influence and shape personality-driven responses and behaviours. This approach enhances personalised communication strategies and fosters deeper, more meaningful connections with diverse audiences.
@@ A.2.1 Model Selection
In this paper, RoBERTa (Robustly optimized BERT approach) Liu et al. [2019] was utilised as the base model. We decided to use RoBERTa after evaluating it by comparing with fine-tuned BERT-base-uncased Devlin [2018] and ALBERT models Chiang et al. [2020]. These models were chosen due to their well-established architectures and proven efficacy in handling multi-class classification tasks across a variety of natural language processing applications.
@@
ALBERT 0.907 0.906781 0.907525 0.907
BERT-base-uncased 0.906 0.906241 0.908288 0.906
RoBERTa (Proposed) 0.919 0.919 0.919 0.919
With an accuracy, F1 score, precision, and recall all at 0.919, RoBERTa demonstrates a consistent and balanced ability to correctly classify instances across all classes. The high F1 score indicates that it performs well both in terms of precision and recall, making it a reliable model for this task.
@@ A.2.2 Training
Following the evaluation done in A.2.1, pretrained RoBERTa model specifically, the RobertaForSequenceClassification model was employed and fine-tuned using the Trainer class provided by the Hugging Face transformers library. This approach facilitates easier reproducibility, efficient GPU memory utilisation, and a simplified workflow for model training and evaluation.
@@ A.2.3 Performance Metrics
  • Weighted Accuracy - Weighted accuracy is the proportion of true results (both true positives and true negatives) among the total number of cases examined, adjusted by the weights of the classes. It measures how often the model is correct overall, taking class imbalance into account. The formula for weighted accuracy is:
    Weighted $\text{Weighted Accuracy}=\sum_{c\in\{\text{Classes}\}}W_{c}\cdot\frac{TP_{c}+TN_{c}}{TP_{c}+TN_{c}+FP_{c}+FN_{c}}$
    where $TP_{c}$ is true positives, $TN_{c}$ is true negatives, $FP_{c}$ is false positives, $FN_{c}$ is false negatives for class $c$, and $W_{c}$ is the weight for class $c$.
  • Weighted F1 Score - The weighted F1 score is the harmonic mean of precision and recall for each class, weighted by the class proportions. It provides a balance between precision and recall across all classes. The formula for the weighted F1 score is:
@@
  • Weighted Precision - Weighted precision is the proportion of true positive results in the predicted positives, adjusted by the weights of the classes. It indicates how many of the predicted positive instances are actually positive, considering class imbalance. The formula for weighted precision is:
@@
  • Weighted Recall - Weighted recall is the proportion of true positive results in the actual positives, adjusted by the weights of the classes. It measures the model’s ability to identify all relevant instances across all classes. The formula for weighted recall is:

### v3

@@ 4.1 Personality Manipulation
4.2 Emoji Generation as a Latent Behaviour
An intriguing behaviour emerged during inference: Mistral-7B-Instruct and LLaMA-2-7B-Chat spontaneously generated emojis in their responses after undergoing PEFT. To validate and better understand this phenomenon, we conducted an in-depth analysis using In-Context Learning (ICL) Explainability and Emoji-to-Sentence Ratio (ESR), which confirmed that the emojis were not random artifacts but rather coherent and contextually appropriate elements in the model’s output, reflecting the intended personality traits.
4.2.1 ESR and ICL Explainability
We verified that the emojis generated were intentional and not random artifacts using ICL explainability. Table 1 shows the results from ICL interpretability and ESR, highlighting frequently used emojis and their corresponding ESRs. Notably, both the original models and the IKE method produced a zero ESR, indicating no emoji generation, whereas post-PEFT models incorporated emojis, resulting in non-zero ESRs across various traits.
Model Method Trait ESR ICL (Emoji)
Mistral-7B-Instruct Original All Traits 0 -
@@ 4.2.1 ESR and ICL Explainability
LLaMA-2-7B-Chat Original All Traits 0 -
@@
LLaMA-3-8B-Instruct Original All Traits 0 -
IKE All Traits 0 -
PEFT All Traits 0 -
Table 1: ESR and ICL Emoji in Generated Text for Different Personality Traits.
@@
As observed in Table 1, LLaMA-2-7B-Chat predominantly produced emojis for Extraversion and Neuroticism, with Extraversion showing the highest emoji-to-sentence ratio at 0.995, where nearly every sentence included an emoji. ICL emojis for Extraversion were positive, such as [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image], reflecting Extraversion qualities. Neuroticism had a ratio of 0.255, using more negative emojis like [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]. Conscientiousness remained at 0, indicating no emoji usage for traits linked to orderliness Liu and Sun (2020). Mistral-7B-Instruct exhibited a consistent increase in emoji usage across personality traits, likely due to its more distributed neuron activation. This was especially pronounced for Openness (0.925) and Conscientiousness (0.82), where creative and productive emojis like [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image] were frequently used. Neuroticism had a moderate ratio (0.575), featuring negative emojis such as [Uncaptioned image] and [Uncaptioned image]. Additionally, we conducted a human evaluation in which participants were asked to assess the relevance of the generated emojis in relation to the target personality traits. Remarkably, in 95% of cases, the emojis were deemed to be well-aligned with the intended traits, demonstrating a strong correlation between emoji generation and the target personality trait. Thus, the PEFT results align with personality traits, enhancing the LLMs’ expressive and contextually appropriate content Kennison et al. (2024).
4.3 Mechanistic Intepretability
Our findings suggest that this emoji generation is likely a result of pre-training on diverse corpora containing emoji patterns Radford et al. (2019), which were subsequently amplified by PEFT. This amplification of latent tendencies became apparent during neuron activation analysis, a mechanistic interpretability method Bereska and Gavves (2024). Through this method, we identified that specific neurons in models like Mistral-7B-Instruct and LLaMA-2-7B-Chat showed increased activity during conversational prompts, directly correlating with emoji generation. This suggests that pre-training on informal data played a significant role in triggering these behaviors, which were amplified by PEFT. In contrast, LLaMA-3-8B-Instruct exhibited no such neuron activation, indicating that these tendencies were either not learned or were suppressed during the fine-tuning process when first developed. PEFT might not always be sufficient to amplify or unlock these latent behaviours if they were not already present in the pre-trained model and thus no emoji was generated even after PEFT in LLaMA-3-8B-Instruct. PEFT, by its nature, makes subtle, localised modifications to the model weights, which could amplify latent behaviours in LLMs Dettmers et al. (2023a). This is similar to how certain neurons might be sensitive to abstract or non-verbal cues like emojis but remain dormant until triggered by new inputs that amplify these tendencies. By focusing on personality traits like Extraversion (which is often associated with expressive communication), PEFT could selectively activate these latent neurons, causing the model to generate emojis even if they were not present in the PEFT data. Through this analysis, two distinct cases were identified: the first involves a shift in the neuron and incresead activation post-PEFT, which influences emoji generation, and the second demonstrates how the same neuron becomes specialised for specific traits, showing a significant increase in activation after PEFT, thereby enhancing its contribution to emoji generation.
4.3.1 Shift in Neuron and Increased Activation After PEFT
Figure 5: Neuron Activation Analysis of original Mistral-7B-Instruct model and PEFT-tuned Mistral-7B-Instruct model for neutral prompt for different Big Five personality traits. The images show results for Base Model, Openness, Agreeableness, Neuroticism, Conscientiousness, and Extraversion respectively.
Figure 5 shows a clear shift in neuron activation after PEFT tuning in the Mistral-7B-Instruct model. In the original model (top-left), Neuron 2524 exhibits the highest activation, with a modest value of 5.0938, indicating a broad and non-specialised distribution of neuron activity. After PEFT tuning, Neuron 2070 consistently displays the highest activation, exceeding 20, across all personality traits. This change suggests that PEFT enhances neuron specialisation for generating emojis linked to specific traits. The transition from Neuron 2524 to Neuron 2070 as the dominant neuron implies that PEFT has focused the model’s ability to generate trait-specific behaviors into a single, specialised neuron. Furthermore, Neuron 2070 consistently exhibited the highest activation across prompts featuring different personality-related emojis and text (Appendix A.6 and A.8). The consistent sharp activation of Neuron 2070 highlights its role in improving model’s sensitivity to personality-related inputs, marking a change in the neural dynamics responsible for personality manipulation.
4.3.2 Increased Activation of the Same Neuron After PEFT
Figure 6: Neuron Activation Analysis of original Llama-2-7B-Chat model and PEFT-tuned Llama-2-7B-Chat model for neutral prompt for different Big Five personality traits. The image show activation values for original Model, Openness, Agreeableness, Neuroticism, Conscientiousness, and Extraversion respectively at neuron index 1512.
In LLaMA-2-7B-Chat (Appendix A.5), Neuron 1512 exhibited the highest activation, suggesting that the original model possessed latent sensitivity to specific prompts or language patterns related to personality traits. Figure 6 and Appendix A.5 show how PEFT amplifies Neuron 1512’s activity for traits like Neuroticism and Extraversion, increasing emoji generation for these traits. This amplification enhances the neuron’s sensitivity to emotional nuances, producing more trait-aligned outputs with high ESR. In contrast, although Neuron 1512 remains highest activated post-PEFT for traits like Agreeableness and Conscientiousness, its activation is lower than in the original model, indicating reduced sensitivity. These traits, being more complex, may require a distributed activation pattern, explaining the limited or inconsistent emoji generation, especially for Conscientiousness (no emojis) and Agreeableness (poorer Trait Alignment scores and low ESR). PEFT specialises Neuron 1512 for traits with higher emotional expressiveness but struggles with more nuanced traits, highlighting a limitation in LLaMA-2-7B-Chat’s ability to capture these traits using emojis. Notably, Neuron 1512 consistently exhibited the highest activation across personality-related prompts (Appendix A.6 and A.8).
5 Limitations
In this paper, we explored the manipulation of personality traits within LLMs, guided by the well-established Big Five personality model. Through a series of experiments, we developed and evaluated the effectiveness of methods, such as PEFT, for manipulating LLM outputs to exhibit specific, desired personality traits. Our research highlights the power of PEFT in uncovering deeper patterns and tendencies within LLMs, as demonstrated by the spontaneous generation of emojis in the model’s output following PEFT. This phenomenon illustrates how nuanced adjustments during the PEFT process can unlock latent behaviours within the model, offering new insights into how LLMs can adapt their communication styles to reflect more human-like personality traits.
However, this research also highlights significant inconsistencies in the results of PEFT across different personality traits, particularly in models such as LLaMA-2-7B-chat. For instance, while the model achieved a high Trait Alignment of 0.975 for neuroticism, it only reached 0.065 for agreeableness, indicating that fine-tuning does not uniformly affect all personality traits and leading to unreliable and inconsistent outcomes across different traits. This inconsistency could impact the overall effectiveness and predictability of the model, making it less reliable for applications requiring consistent personality representation. Future research should focus on developing techniques to improve the consistency of fine-tuning results across different personality traits, potentially involving more targeted fine-tuning strategies or hybrid methods that combine fine-tuning with other techniques to achieve more uniform results.
Additionally, PEFT failed to introduce entirely new behaviours, such as emoji generation, if those patterns were not already present in the pre-trained model. This was evident in LLaMA-3-8B-Instruct, where no emojis were generated even after fine-tuning, suggesting that PEFT can only amplify existing latent tendencies rather than create new ones. Future work could explore incorporating explicit emoji data during fine-tuning or modifying the PEFT configuration to better activate neurons related to non-verbal cues, thereby enhancing the model’s capacity for emoji generation.
Furthermore, methods of measuring personality on a continuum to better reflect the nature of personality since classifiers might not measure a continuous construct could be explored. Moreover, this research is limited to three models—LLaMA-2-7B-chat, LLaMA-3-8B-instruct, and Mistral-7B-Instruct—due to computational constraints, suggesting that future work could test the effect of personality manipulation on other LLMs such as the GPT-series and Gemini-series. Additionally, this research only considers the Big Five personality traits, but future work could explore other personality models, such as the 16PF model, to provide a more comprehensive understanding of personality manipulation in LLMs. Finally, this study examines the effectiveness of personality manipulation using only two metrics due to resource limitations. Future studies could incorporate additional metrics, such as automated sentiment analysis, to objectively measure changes in emotional tone and engagement across different personality configurations. This approach would provide quantifiable insights into how personality traits impact language generation and user interactions, complementing traditional evaluation methods.
6 Ethical Considerations
The integration of personality traits and emoji generation into large language models (LLMs) using Parameter-Efficient Fine-Tuning (PEFT) introduces several ethical concerns that must be addressed to ensure responsible deployment.
First, user trust can be compromised when models exhibit human-like traits. Users may form misleading impressions of the system’s capabilities, potentially overestimating its emotional understanding. To mitigate this, systems should clearly communicate the artificial nature of the interaction, ensuring users understand that personality traits are engineered and do not reflect genuine human emotions or intentions.
Second, bias amplification poses a significant risk. Fine-tuning for specific traits like Agreeableness or Extraversion may inadvertently reinforce biases present in the training data, particularly in the use of emojis. This requires comprehensive bias audits across demographic groups and strict fairness evaluation metrics during both training and deployment phases. In domains like customer service or mental health, the use of LLMs can lead to emotional manipulation through personality-driven responses. Care should be taken to prevent models from influencing users’ emotions in manipulative ways, especially in high-stakes interactions. Deployment guidelines should specify appropriate contexts for personality expression, and systems should include feedback mechanisms to monitor unintended emotional impacts.
Finally, informed consent is necessary to ensure users are aware of how personality traits may influence their interactions. Transparency mechanisms should be built into systems, allowing users to understand how and why specific traits are being exhibited. Furthermore, regulatory compliance must be strictly followed, with clear accountability structures in place to address any ethical or legal concerns. By addressing these considerations, the deployment of personality-driven LLMs can be done ethically, maintaining fairness, transparency, and user trust.
Appendix A Appendix
@@ A.1 Personality Manipulation Dataset
The Personality Manipulation dataset consists of $5000$ instances, with $4000$ allocated for training and $1000$ for testing. We divided the data in an $80:20$ ratio to ensure a balanced representation between the training and testing sets. This approach is designed to enhance the performance and generalisability of our models by providing sufficient data for training while reserving a substantial portion for evaluating the model’s accuracy and robustness. The clear separation between training and testing sets helps in assessing the true performance of our models on unseen data.
@@
Instruction: Exhibit the trait of Target Personality when answering the question to express opinion on the certain Edit Topic.
@@
Table 2: Prompt used for Few-shot Prompting in Dataset Generation.
@@
For instance, as visible in the figure 8, the word cloud for extraversion showcases words that reflect the energetic, outgoing, and sociable nature of individuals with this personality trait. Prominent terms such as "absolutely," "always," "believe," "love," and "thrilling" emphasise the enthusiasm, positivity, and social engagement typical of extraverts. Similarly, the word cloud for neuroticism reveals words associated with emotional instability, anxiety, and self-consciousness. Key terms like "feel," "worry," "sometimes," "make," and "bit" are prominently displayed, reflecting the frequent emotional fluctuations and concerns typical of neurotic individuals. Words like "anxious," "overwhelming," and "struggle" further emphasise the challenges faced by those with high levels of neuroticism. Collectively, these word clouds provide a rich visual representation of the lexical diversity associated with each personality trait, offering valuable insights into the distinct characteristics and behavioral tendencies that define them.
@@ A.1.1 Text Analysis
In conclusion, LDA-based topic modeling reveals significant insights into how different themes resonate with various personality traits. This understanding is crucial for the manipulation of personalities in language models, allowing for the creation of content that is tailored to engage specific personality profiles effectively. By leveraging these insights, content can be crafted to not only align with individual preferences but also to influence and shape personality-driven responses and behaviors. This approach enhances personalised communication strategies and fosters deeper, more meaningful connections with diverse audiences.
@@ A.2.1 Model Selection
RoBERTa (Robustly optimized BERT approach) Liu et al. (2019) was utilised as the base model in this study. We decided to use RoBERTa after evaluating it by comparing with fine-tuned BERT-base-uncased Devlin (2018) and ALBERT models Chiang et al. (2020). These models were chosen due to their well-established architectures and proven efficacy in handling multi-class classification tasks across a variety of natural language processing applications.
@@
ALBERT 0.907 0.9068 0.9075 0.907
BERT-base-uncased 0.906 0.9062 0.9083 0.906
RoBERTa
(Proposed) 0.919 0.919 0.919 0.919
With an accuracy, F1 score, precision, and recall value of 0.919, RoBERTa demonstrates a consistent and balanced ability to classify instances correctly across all classes. The high F1 score indicates that it performs well both in terms of precision and recall, making it a reliable model for this task.
@@ A.2.2 Training
Following the evaluation described in A.2.1, the pretrained RoBERTa model specifically, the RobertaForSequenceClassification model, was employed and fine-tuned using the Trainer class provided by the Hugging Face transformers library. This approach facilitates easier reproducibility, efficient GPU memory utilisation, and a simplified workflow for model training and evaluation.
@@ A.2.3 Performance Metrics
  • Weighted Accuracy (WA) - The proportion of true results (both true positives and true negatives) among the total number of cases, adjusted by class weights. It measures overall correctness, accounting for class imbalance:
    $\text{WA}=\sum_{c\in\{\text{Classes}\}}W_{c}\cdot\frac{TP_{c}+TN_{c}}{TP_{c}+TN_{c}+FP_{c}+FN_{c}}$
  • Weighted F1 Score - The harmonic mean of precision and recall for each class, weighted by class proportions. It balances precision and recall:
@@
  • Weighted Precision - The proportion of true positive results in predicted positives, adjusted by class weights. Indicates the accuracy of positive predictions:
@@
  • Weighted Recall - The proportion of true positive results in actual positives, adjusted by class weights. Measures the model’s ability to identify relevant instances: